7. A computer is infected with the Sasser virus. Assume that it infects 20 other computers within 5 minutes; and that these PCs and servers each infect 20 more machines within another five minutes, etc. How long until 100 million computers are infected?
I am dumbfounded by this and need explanation. I need to see the work here. thank you.

Answer by stanbon(48545)

A computer is infected with the Sasser virus. Assume that it infects 20 other computers within 5 minutes; and that these PCs servers each infect 20 more machines within another five minutes, etc. How long until 100 million computers are infected?
---------------------
When you have a problem like this you should construct a table of pairs of
points; then try to see a pattern.
0,1
That 1 gets multiplied by 20 after 5 seconds
------
5,1+20
That 1+20 gets multiplied by 20 after another 5 seconds
-------------------------
10,(1+20)*20
Muliplied by 20 again
---------------------------------
15,[(1+20)*20]*20 = [(1+20)]*20^2
20,[(1+20)]*20^3
25,[(1+20)]*20^(25/5)-1)
30,[(1+20)]*20^((30/5)-1)
...
t seconds , [(1+20)]*20^((t/5)-1)
--------------------------------------------
Equation:
A(t) = (1+20)*20^((t/5)-1)
--------------------------------
How long until 100 million computers are infected?
Solve for t in seconds: (1+20)*20^((t/5)-1) = 100 million
20^((t/5)-1)] = (100/21) million
Take the log of both sides to get:
((t/5)-1)log(20) = log(100 million) - log(21)
((t/5)-1)log(20) = 8 - 1.3222 = 6.6778
[(t/5)-1] = 6.6778/1.3010 = 5.1327
(t/5) = 6.1327
t = 30.6635 seconds
or
time = 30.6635/60 = 0.5111 minutes
====================================
Comment: I checked this out. You actually get 100,003,673 infected computers.
The additional 3673 computers is there because we are rounding off the
numbers as we do the calculations.
=====================================
Cheers,
Stan H.